A teacher purchased 20 calculators and 10 measuring tapes for her class and paid $495. Later, she realized that she didn't order enough supplies. She placed another order of 8 of the same calculators and 1 more of the same measuring tape and paid $178.90. How much does one calculator cost? How much does one measuring tape cost?
1
Expert's answer
2020-04-27T18:26:45-0400

Let "x" be a price of one calculator and "y" be a price of one measuring tape.

A teacher paid $495 for 20 calculators and 10 measuring tapes


"20x+10y=495"

A teacher paid $178.90 for 8 calculators and 1 measuring tape


"8x+y=178.90"

"2x+y=49.50""8x+y=178.90"

"6x=129.40""y=49.50-2x"

"x={64.70 \over 3}""y=49.50-2x"

"\dfrac{64,70}{3}" is not integer. I think the condition is wrong.

She placed another order of 8 of the same calculators and 1 more of the same measuring tape and paid $178.50.


"20x+10y=495""8x+y=178.50"

"2x+y=49.50""8x+y=178.50"

"6x=129""y=49.50-2x"


"x=21.50""y=6.50"

One measuring tape costs $6.50